Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

If the probability of Sita getting selected to a school...

Expert replies
by BTGmoderatorLU » Sun Jan 21, 2018 7:59 am
If the probability of Sita getting selected to a school is 20% and she applied to 3 schools. What is the probability that she will get selected in at least one school?

A. 24/125
B. 21/125
C. 61/125
D. 12/125
E. 18/125

The OA is C.

I'm confused with this PS question. Experts, any suggestion about how to solve? I don't understand this question. Thanks in advance.
Join the discussion
Source: — Problem Solving |

by [email protected] » Sun Jan 21, 2018 10:50 am
Hi LUANDATO,

We're told that the probability of Sita getting selected to any individual school is 20% and that she applied to 3 schools. We're asked for the probability that she will get selected to AT LEAST one school.

With these types of probability questions, it's often easiest to calculate the probability of what you DON'T WANT - and then subtract that fraction from 1 to determine the probability of what you DO want. The question asks for the probability that she will get selected to AT LEAST one school, so we'll calculate the probability that she will get selected to 0 schools. A 20% chance of getting selected means an 80% chance of NOT getting selected. Since the answers are fractions, we should write 80% as 4/5:

(Not selected)(Not selected)(Not selected) = (4/5)(4/5)(4/5) = 64/125 chance of getting selected by 0 schools

Thus, he probability that she will get selected to AT LEAST one school is... 1 - 64/125 = 61/125

Final Answer: C

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by EconomistGMATTutor » Sun Jan 21, 2018 11:47 am
If the probability of Sita getting selected to a school is 20% and she applied to 3 schools. What is the probability that she will get selected in at least one school?

A. 24/125
B. 21/125
C. 61/125
D. 12/125
E. 18/125

The OA is C.

I'm confused with this PS question. Experts, any suggestion about how to solve? I don't understand this question. Thanks in advance.
Hi LUANDATO,
Let's take a look at your question.

The probability of Sita getting selected to a school = 20% = 20/100 = 1/5
This question can easily be solved by calculating the probability of opposite event.
The probability of Sita not getting selected to a school = 1 - 1/5 = 4/5

Probability that Sita not getting selected at all three schools can be calculated as:
$$=\left(\frac{4}{5}\right)^3=\frac{65}{125}$$

Now we can solve the probability of Sita getting selected in atleast one school as:
$$=1-\frac{64}{125}=\frac{125-64}{125}=\frac{61}{125}$$

Therefore, Option C is correct.

Hope it helps.
I am available if you's like any follow up.
GMAT Prep From The Economist
We offer 70+ point score improvement money back guarantee.
Our average student improves 98 points.

Image
Join the discussion

by Jeff@TargetTestPrep » Tue Jan 23, 2018 9:11 am
LUANDATO wrote:If the probability of Sita getting selected to a school is 20% and she applied to 3 schools. What is the probability that she will get selected in at least one school?

A. 24/125
B. 21/125
C. 61/125
D. 12/125
E. 18/125
We can use the equation:

P(selected into at least 1 school) = 1 - P(selected to no schools)

P(selected to no schools) = 4/5 x 4/5 x 4/5 = 64/125

Thus,

P(selected into at least 1 school) = 1 - 64/125 = 61/125

Answer: C

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion